# Effective algorithms for tensor train decomposition via the UTV framework

### v1

@@ III-A TT-ULV case
0:  $d$-dimensional tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0:  Left-orthogonal TT-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
@@
4:   Compute the truncated rank-$r_{k}$ approximation from ULV decomposition $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{E}$, where $\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$ with $\widehat{\mathbf{L}}\in\mathbb{R}^{r_{k}\times r_{k}}$ is the leading term and $\mathbf{E}$ denotes the residual part in the ULV decomposition (5).
@@
6:   $\mathbf{C}=\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$.
@@
8:  $\mathcal{G}^{(d)}=\mathbf{C}$.
To facilitate the presentation of error analysis, we introduce some notations. The TT-ULV algorithm 1 adopts the left-to-right sweep (steps 2-7). In the $k$th iteration, step 3 computes a temporary matrix $\mathbf{C}\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\cdots I_{d}}$, we denote its corresponding $(d-k+2)$th-order tensor by $\mathcal{A}^{(k)}=\mathrm{reshape}(\mathbf{C},[r_{k-1},I_{k},I_{k+1},\cdots,I_{d}])\in\mathbb{R}^{r_{k-1}\times I_{k}\times\cdots\times I_{d}}$, and the $(d-k+1)$th-order tensor by $\widetilde{\mathcal{A}}^{(k)}=\mathrm{reshape}(\mathbf{C},[r_{k-1}I_{k},I_{k+1},\cdots,I_{d}])\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\times\cdots\times I_{d}}$. Step 4 computes the column-orthogonal matrix $\widehat{\mathbf{U}}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}}$, denoted by $\mathbf{U}^{(k)}$ in the $k$th iteration. The whole computational processes generate temporary tensors $\{\mathcal{A}^{(k)}\}_{k=1}^{d-1}$, $\{\widetilde{\mathcal{A}}^{(k)}\}_{k=1}^{d-1}$, and $d-1$ column-orthogonal matrices $\{\mathbf{U}^{(k)}\}_{k=1}^{d-1}$ that give birth to the TT-cores. We firstly derive the following proposition and lemma.
Proposition III.1.
@@
for $l=1,\dots,d-k+1$ for all $k=1,\cdots,d-1$, where $\mathbf{A}_{l}^{(k)}$ is the $l$th unfolding matrix of tensor $\mathcal{A}^{(k)}\in\mathbb{R}^{r_{k-1}\times I_{k}\times\cdots\times I_{d}}$, $\mathbf{I}_{m}\in\mathbb{R}^{m\times m}$ is the identity matrix, and $\otimes$ denotes the matrix Kronecker product.
@@
where $\mathbf{A}_{2}^{(k)}\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\cdots I_{d}},\mathbf{U}^{(k)}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}},\mathbf{A}_{1}^{(k+1)}\in\mathbb{R}^{r_{k}\times I_{k+1}\cdots I_{d}}$. From the orthogonality $\mathbf{U}^{(k)\top}\mathbf{E}=\mathbf{O}$,
@@
From the definitions of the notations $\mathcal{A}^{(k)}$ and $\widetilde{\mathcal{A}}^{(k)}$, the $l$th unfolding matrices of $\widetilde{\mathcal{A}}^{(k)}$ are exactly the $(l+1)$th unfolding matrix of $\mathcal{A}^{(k)}$, i.e., $\mathbf{A}^{(k)}_{l+1}$, for $l=1,\cdots,d-k+1$. Clearly, the above equation (8) can be rewritten as the tensor format
@@
From the above tensor equation, the equations (7) hold by Proposition 2.1. ∎
Lemma III.1.
Let $m\geq s\geq r$ be positive integers, $\mathbf{Q}\in\mathbb{R}^{s\times m}$ with $\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{s}$, $\mathbf{A}\in\mathbb{R}^{m\times n}$, and $\mathbf{B}=\mathbf{Q}\mathbf{A}\in\mathbb{R}^{s\times n}$. Then the rank-r truncated error of the ULV (or URV) decomposition of $\mathbf{B}$ can be smaller than that of $\mathbf{A}$.
@@
Since $\mathbf{Q}$ has orthogonal rows with unit norm, $\|\mathbf{Q}\|_{2}=1$, $\|\mathbf{Q}\mathbf{E}\|_{F}\leq\|\mathbf{E}\|_{F}$. For $\mathbf{Q}\widehat{\mathbf{U}}\in\mathbb{R}^{s\times r}$, we can perform the Gram-Schmidt orthogonalization process to its columns to obtain a decomposition (similar to the QR decomposition) $\mathbf{Q}\widehat{\mathbf{U}}=\widetilde{\mathbf{U}}\mathbf{L}$, where $\widetilde{\mathbf{U}}\in\mathbb{R}^{s\times r}$ has orthogonal columns with unit norm and $\mathbf{L}\in\mathbb{R}^{r\times r}$ is lower-triangular. Then $\widetilde{\mathbf{L}}=\mathbf{L}\widehat{\mathbf{L}}$ is still a lower-triangular matrix. Therefore,
@@
the leading term $\widetilde{\mathbf{U}}\widetilde{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$ is actually a rank-$r$ truncated ULV decomposition of $\mathbf{B}$. Thus, it finds a rank-$r$ truncated ULV decomposition for $\mathbf{B}$ such that the approximate error $\|\mathbf{Q}\mathbf{E}\|_{F}$ is smaller than that of $\mathbf{A}$. For the URV case, the above analysis similarly holds. ∎
Remark III.1.
From the properties of the Kronecker product, it is easily verified that the Kronecker matrices $\left(\mathbf{I}_{I_{k+l-1}}\otimes\cdots\otimes\mathbf{I}_{I_{k+1}}\otimes\mathbf{U}^{(k)\top}\right)$ in the equation (7) are row-orthogonal with unit-norm rows.
Next, based on Proposition III.1 and Lemma III.1 the error analysis for the TT-ULV algorithm 1 can be derived, which can be seen as the generalizations of the results for TT-SVD [29].
Theorem III.1 (TT-ULV case).
Given a tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Assume that the rank-$r_{k}$ approximation error of its $k$th unfolding matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ from the ULV decomposition (5) is $\varepsilon_{k}$, that is,
$\mathbf{A}_{k}=\mathbf{U}^{(k)}_{1}\mathbf{L}^{(k)}_{11}\mathbf{V}^{(k)\top}_{1}+\mathbf{E}^{(k)},$
where $\mathbf{L}^{(k)}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$, $\mathbf{E}^{(k)}=\mathbf{U}^{(k)}_{2}(\mathbf{L}^{(k)}_{21}\mathbf{V}^{(k)\top}_{1}+\mathbf{L}^{(k)}_{22}\mathbf{V}^{(k)\top}_{2})$ and $\|\mathbf{E}^{(k)}\|_{F}=\varepsilon_{k}$ for all $k=1,\dots,d-1$. Then the approximate tensor $\widehat{\mathcal{A}}$ with TT-ranks not higher than $r_{k}$’s, computed via TT-ULV algorithm 1 in the left-to-right sweep, meets
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$ (9)
@@
It yields the first TT-core from the ULV decomposition of the first unfolding matrix
$\mathbf{A}_{1}=\mathbf{U}_{1}^{(1)}\mathbf{A}_{2}^{(2)}+\mathbf{E}^{(1)},$ (10)
where $\mathbf{U}^{(1)}_{1}\in\mathbb{R}^{I_{1}\times r_{1}}$ leads to the first TT-core, and $\mathbf{A}_{2}^{(2)}=\mathbf{L}^{(1)}_{11}\mathbf{V}^{(1)\top}_{1}\in\mathbb{R}^{r_{1}\times(I_{2}\cdots I_{d})}$ is used to produce the subsequent TT-cores. Based on the column orthogonality, it follows that $\mathbf{U}_{1}^{(1)\top}\mathbf{E}^{(1)}=\mathbf{O}$. As we can see, the subsequent processes are to approximate the $(d-1)$th-order tensor $\widetilde{\mathcal{A}}^{(2)}=\mathrm{reshape}(\mathbf{A}_{2}^{(2)},[r_{1}I_{2},I_{3},\cdots,I_{d}])$ folded from the matrix $\mathbf{A}_{2}^{(2)}$. We denote by $\mathcal{B}\in\mathbb{R}^{r_{1}I_{2}\times I_{3}\times\cdots\times I_{d}}$ the approximation of $\widetilde{\mathcal{A}}^{(2)}$ computed by the algorithm, and $\mathbf{B}=\mathrm{reshape}(\mathcal{B},[r_{1},(I_{2}\cdots I_{d})])$. Then the approximate error of the output tensor $\widehat{\mathcal{A}}$ of the TT-ULV algorithm is
@@
where the third inequality is from the orthogonality of $\mathbf{U}_{1}^{(1)\top}\mathbf{E}^{(1)}=\mathbf{O}$ and $\|\mathbf{U}_{1}^{(1)}\|_{2}=1$. Therefore, it reduces to the estimation of the approximate error for the $(d-1)$th-order tensor $\widetilde{\mathcal{A}}^{(2)}$ by TT-ULV algorithm. As explained in Proposition III.1, Lemma III.1 and Remark III.1, the the truncated error of rank-$r_{k}$ approximation from the ULV decomposition of the $(k-1)$th unfolding matrix of $\widetilde{\mathcal{A}}^{(2)}$ can be smaller than the approximate error $\varepsilon_{k}$ of the rank-$r_{k}$ truncated ULV decomposition for $\mathbf{A}_{k}$, for all $k=1,\dots,d-1$. Hence, Based on the above analysis and mathematical induction method, we get
@@
Combine it with the equation (11), the theorem is proved. ∎
From the analysis of Theorem III.1, we can also bound the total approximate error of the TT-UTV algorithm 1 by the local errors of each step in the truncated UTV decompositions.
Corollary III.1.
For the TT-ULV algorithm 1 in the left-to-right sweep, if the error of the $k$th truncated ULV decomposition in step 4 is $\delta_{k}$, then the total error of the computed TT-approximation tensor $\widehat{\mathcal{A}}$ satisfies $\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\delta_{k}^{2}}$.
Therefore, the version of the TT-ULV algorithm in the left-to-right sweep for prescribed accuracy can be easily formulated by referring to Algorithm 1 and the TT-SVD [29, Algorithm 1].
@@
0:  A $d$-dimensional tensor $\mathcal{A}$, prescribed relative error $\varepsilon$.
0:  Left-orthogonal TT-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation tensor $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$ not higher than the ULV truncated ranks of the unfolding matrices $\mathbf{A}_{k}$ such that the error $\delta=\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$. Then the computed approximation satisfies
@@
1:  Compute truncation parameter $\delta=\frac{\varepsilon}{\sqrt{d-1}}\|\mathcal{A}\|_{F}$, set $r_{0}=1$.
@@
5:   Compute $\delta$-truncated ULV decomposition $\mathbf{C}=\widehat{\mathbf{U}}\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}+\mathbf{E}$, $\|\mathbf{E}\|_{F}\leq\delta$, $\widehat{\mathbf{L}}\in\mathbb{R}^{r_{k}\times r_{k}}$.
@@
7:   $\mathbf{C}:=\widehat{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$.
@@
9:  $\mathcal{G}^{(d)}=\mathbf{C}$.
An essential requisite for the above error analysis for Algorithm 1 is that in each iteration the column-orthogonal matrices $\widehat{\mathbf{U}}$ used to generate the TT-cores are orthogonal to the residual part $\mathbf{E}$. Then the accumulation of the truncated errors can be theoretically controlled by a sharp bound (9).
However, if the right-orthogonal TT-cores are needed, the TT-ULV algorithm is theoretically not recommended to be employed in the right-to-left sweep, especially for the computations of the approximate TT tensor under a prescribed accuracy. Since the column-orthogonal matrix $\widehat{\mathbf{V}}$ in truncated ULV decomposition is not orthogonal to $\mathbf{E}$, the above analysis for approximate error is no longer valid, and only a larger error bound can be obtained. More details can be seen in the APPENDIX. Interestingly, the URV case is precisely the opposite.
III-B TT-URV case
For the URV decomposition (3), the matrix $\mathbf{V}_{1}$ is orthogonal to the residual part $(\mathbf{U}_{1}\mathbf{R}_{12}+\mathbf{U}_{2}\mathbf{R}_{22})\mathbf{V}_{2}^{\top}$. Wherefore, using the URV decomposition in the right-to-left sweep to compute TT format is more suitable. Algorithm 3 presents the TT-URV algorithm for fixed TT-ranks in the right-to-left sweep to compute the TT-format tensor with right-orthogonal cores.
The results similar to Proposition III.1 and Theorem III.1 also hold for Algorithm 3.

### v2

@@ 3.1 TT-ULV case
0: A $d$th-order tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$, and fixed TT-ranks $\{r_{0},r_{1},\dots,r_{d}\}$ with $r_{0}=r_{d}=1$.
0: Left-orthogonal $T$-cores $\{\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}\}$ of the approximation $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$’s.
@@
4:  Compute the truncated rank-$r_{k}$ approximation from ULV decomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}_{1}^{\top}+\mathbf{E}$, where $\mathbf{L}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$ is the leading term of $\mathbf{L}$, and $\mathbf{E}$ denotes the residual part in Eq. (4).
@@
6:  $\mathbf{C}=\mathbf{L}_{11}\mathbf{V}_{1}^{\top}$.
@@
[…]
To facilitate the presentation of error analysis, we introduce some notations. The TT-ULV algorithm 1 adopts the left-to-right sweep (steps 2-7). In the $k$th iteration, step 3 computes a temporary matrix $\mathbf{C}\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\cdots I_{d}}$; we denote its corresponding (dk+2)(d-k+2)th-order tensor by $\mathcal{A}^{(k)}=\mathrm{reshape}(\mathbf{C},[r_{k-1},I_{k},I_{k+1},\cdots,I_{d}])\in\mathbb{R}^{r_{k-1}\times I_{k}\times\cdots\times I_{d}}$, and the (dk+1)(d-k+1)th-order tensor by $\widetilde{\mathcal{A}}^{(k)}=\mathrm{reshape}($ $\mathbf{C},[r_{k-1}I_{k},I_{k+1},\cdots,I_{d}])\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\times\cdots\times I_{d}}$. Step 4 computes the column-orthogonal matrix $\mathbf{U}_{1}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}}$, denoted by $\mathbf{U}^{(k)}$ in the $k$th iteration. With the notations introduced, the whole process generates temporary tensors $\{\mathcal{A}^{(k)}\}_{k=1}^{d-1}$, $\{\widetilde{\mathcal{A}}^{(k)}\}_{k=1}^{d-1}$, and d1d-1 column-orthogonal matrices $\{\mathbf{U}^{(k)}\}_{k=1}^{d-1}$ that give rise to the TT-cores.
Proposition 3.1.
@@
for $l=1,\dots,d-k+2$ for all $k=1,\cdots,d-1$, where $\mathbf{A}_{l}^{(k)}$ is the $l$th unfolding matrix of tensor $\mathcal{A}^{(k)}\in\mathbb{R}^{r_{k-1}\times I_{k}\times I_{k+1}\times\cdots\times I_{d}}$.
@@
where $\mathbf{A}_{2}^{(k)}\in\mathbb{R}^{r_{k-1}I_{k}\times I_{k+1}\cdots I_{d}},\mathbf{U}^{(k)}\in\mathbb{R}^{r_{k-1}I_{k}\times r_{k}},\mathbf{A}_{1}^{(k+1)}\in\mathbb{R}^{r_{k}\times I_{k+1}\cdots I_{d}}$. According to the orthogonality $\mathbf{U}^{(k)\top}\mathbf{E}=\mathbf{O}$, we have
@@
From the definitions of the notations $\mathcal{A}^{(k)}$ and $\widetilde{\mathcal{A}}^{(k)}$, the $l$th unfolding matrices of $\widetilde{\mathcal{A}}^{(k)}$ are exactly the $(l+1)$th unfolding matrix of $\mathcal{A}^{(k)}$, i.e., $\mathbf{A}^{(k)}_{l+1}$, for $l=1,\cdots,d-k+1$. Clearly, the above Eq. (9) can be rewritten in tensor format
@@
From the above tensor equation, the Eq. (8) holds by Proposition 2.1. ∎
Lemma 3.1.
Let $m\geq s\geq r$ be positive integers, $\mathbf{Q}\in\mathbb{R}^{s\times m}$ be a row-orthogonal matrix with unit norm rows, $\mathbf{A}\in\mathbb{R}^{m\times n}$, and $\mathbf{B}=\mathbf{Q}\mathbf{A}\in\mathbb{R}^{s\times n}$. Then the rank-r truncated error of the ULV (or URV) decomposition of $\mathbf{B}$ can be smaller than that of $\mathbf{A}$.
@@
Since $\mathbf{Q}$ has orthogonal rows with unit norm, $\|\mathbf{Q}\|_{2}=1$, then $\|\mathbf{Q}\mathbf{E}\|_{F}\leq\|\mathbf{E}\|_{F}=\varepsilon$. For $\mathbf{Q}\widehat{\mathbf{U}}\in\mathbb{R}^{s\times r}$, we can perform the Gram-Schmidt orthogonalization process on its columns to obtain a decomposition (similar to the QR decomposition) $\mathbf{Q}\widehat{\mathbf{U}}=\widetilde{\mathbf{U}}\mathbf{L}$, where $\widetilde{\mathbf{U}}\in\mathbb{R}^{s\times r}$ has orthogonal columns with unit norm and $\mathbf{L}\in\mathbb{R}^{r\times r}$ is lower-triangular. Then $\widetilde{\mathbf{L}}=\mathbf{L}\widehat{\mathbf{L}}$ is still a lower-triangular matrix. Therefore,
@@
the leading term $\widetilde{\mathbf{U}}\widetilde{\mathbf{L}}\widehat{\mathbf{V}}^{\top}$ is actually a rank-$r$ truncated ULV decomposition of $\mathbf{B}$, and the residual part $\|QE\|_{F}\leq\varepsilon$. Thus, it finds a rank-$r$ truncated ULV decomposition for $\mathbf{B}$ such that the approximate error is smaller than that of $\mathbf{A}$. For the URV case, the above analysis similarly holds. ∎
Remark 3.1.
According to Proposition 2.2, the Kronecker matrices $\left(\mathbf{I}_{I_{k+l-1}}\otimes\cdots\otimes\mathbf{I}_{I_{k+1}}\otimes\mathbf{U}^{(k)\top}\right)$ in the Eq. (8) are row-orthogonal with unit-norm rows.
Next, the error analysis for the TT-ULV algorithm 1 can be derived as shown in Theorem 3.1 and Corollary 3.1, which can be seen as a generalization of the results for TT-SVD [oseledets2011tensor].
Theorem 3.1 (TT-ULV case).
Given a tensor $\mathcal{A}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$. Assume that the rank-$r_{k}$ approximation error of its $k$th unfolding matrix $\mathbf{A}_{k}\in\mathbb{R}^{(I_{1}\cdots I_{k})\times(I_{k+1}\cdots I_{d})}$ from the ULV decomposition is $\varepsilon_{k}$, that is,
$\mathbf{A}_{k}=\begin{bmatrix}\mathbf{U}^{(k)}_{1}&\mathbf{U}^{(k)}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{L}^{(k)}_{11}&\mathbf{O}\\ \mathbf{L}^{(k)}_{21}&\mathbf{L}^{(k)}_{22}\end{bmatrix}\begin{bmatrix}\mathbf{V}^{(k)\top}_{1}\\ \mathbf{V}^{(k)\top}_{2}\end{bmatrix}=\mathbf{U}^{(k)}_{1}\mathbf{L}^{(k)}_{11}\mathbf{V}^{(k)\top}_{1}+\mathbf{E}^{(k)},$ (10)
where $\mathbf{L}^{(k)}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$, $\mathbf{E}^{(k)}=\mathbf{U}^{(k)}_{2}(\mathbf{L}^{(k)}_{21}\mathbf{V}^{(k)\top}_{1}+\mathbf{L}^{(k)}_{22}\mathbf{V}^{(k)\top}_{2})$ and $\|\mathbf{E}^{(k)}\|_{F}=\varepsilon_{k}$ for all $k=1,\dots,d-1$. Then the approximate tensor $\widehat{\mathcal{A}}$ with TT-ranks not higher than $\{r_{k}\}_{0}^{d}$, computed via the TT-ULV algorithm 1 in a left-to-right sweep, meets
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$
@@
In the TT-ULV algorithm 1, it yields the first TT-core from the ULV decomposition of the first unfolding matrix
$\mathbf{A}_{1}=\mathbf{U}_{1}^{(1)}\mathbf{A}_{2}^{(2)}+\mathbf{E}^{(1)},$
where $\mathbf{U}^{(1)}_{1}\in\mathbb{R}^{I_{1}\times r_{1}}$ leads to the first TT-core, and $\mathbf{A}_{2}^{(2)}=\mathbf{L}^{(1)}_{11}\mathbf{V}^{(1)\top}_{1}\in\mathbb{R}^{r_{1}\times(I_{2}\cdots I_{d})}$ is used to produce the subsequent TT-cores. Note that $\mathbf{A}_{2}^{(2)}$ is the second unfolding matrix of the tensor $\mathcal{A}^{(2)}$. Based on the column orthogonality, it follows that $\mathbf{U}_{1}^{(1)\top}\mathbf{E}^{(1)}=\mathbf{O}$. As we can see, the subsequent processes are to approximate the (d1)(d-1)th-order tensor $\widetilde{\mathcal{A}}^{(2)}=\mathrm{reshape}(\mathbf{A}_{2}^{(2)},[r_{1}I_{2},I_{3},\cdots,I_{d}])$ folded from the matrix $\mathbf{A}_{2}^{(2)}$. We denote by $\mathcal{B}\in\mathbb{R}^{r_{1}I_{2}\times I_{3}\times\cdots\times I_{d}}$ the approximation of $\widetilde{\mathcal{A}}^{(2)}$ computed by the algorithm, and $\mathbf{B}=\mathrm{reshape}(\mathcal{B},[r_{1},(I_{2}\cdots I_{d})])$. Then, the approximate error of the output tensor $\widehat{\mathcal{A}}$ of the TT-ULV algorithm is
@@
where the third inequality is derived from the orthogonality of $\mathbf{U}_{1}^{(1)\top}\mathbf{E}^{(1)}=\mathbf{O}$ and $\|\mathbf{U}_{1}^{(1)}\|_{2}=1$. Therefore, it reduces to the estimation of the approximate error for the (d1)(d-1)th-order tensor $\widetilde{\mathcal{A}}^{(2)}$ by TT-ULV algorithm. As explained in Proposition 3.1, Lemma 3.1, and Remark 3.1, the truncated error of rank-$r_{k}$ approximation from the ULV decomposition of the (k1)(k-1)th unfolding matrix of $\widetilde{\mathcal{A}}^{(2)}$ can be smaller than the approximate error $\varepsilon_{k}$ of the rank-$r_{k}$ truncated ULV decomposition for $\mathbf{A}_{k}$, for all $k=1,\dots,d-1$. Hence, based on the above analysis and by the method of mathematical induction, we get
@@
Combine it with the Eq. (11), the theorem is proved. ∎
As established by Theorem 3.1, the approximation error of the TT-ULV algorithm is bounded by the low-rank approximation errors of the ULV decompositions applied to its unfolding matrices. This result also reveals that the low-rankness of these unfolding matrices directly implies a low-rank TT structure. Furthermore, the analysis also allows us to obtain the final approximation error of the TT-UTV algorithm 1, showing how it propagates from the local errors incurred at each step of the sequential truncated UTV decompositions.
Corollary 3.1.
For the TT-ULV algorithm 1 in the left-to-right sweep, if the error of the $k$th truncated ULV decomposition in step 4 is $\varepsilon_{k}$, then the total error of the computed TT-approximation tensor $\widehat{\mathcal{A}}$ satisfies
$\|\mathcal{A}-\widehat{\mathcal{A}}\|_{F}\leq\sqrt{\sum_{k=1}^{d-1}\varepsilon_{k}^{2}}.$
@@
0: A $d$th-order tensor $\mathcal{A}$, and prescribed relative error tolerance $\varepsilon$, and the error weights $\{w_{k}\}_{1}^{d-1}$ with $\sum_{k=1}^{d-1}w_{k}^{2}=1$.
0: Left-orthogonal $T$-cores $\mathcal{G}^{(1)},\dots,\mathcal{G}^{(d)}$ of the approximation tensor $\widehat{\mathcal{A}}$ with TT-ranks $r_{k}$ not higher than the ULV truncated ranks of the unfolding matrices $\mathbf{A}_{k}$ such that the truncated error $\varepsilon_{k}=w_{k}\varepsilon\|\mathcal{A}\|_{F}$. Then from Corollary 3.1, the computed TT-approximation tensor satisfies
@@
1: Compute truncation parameters $\varepsilon_{k}=w_{k}\varepsilon\|\mathcal{A}\|_{F}$, set $r_{0}=1$.
@@
5:  Compute truncated ULV decomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}_{1}^{\top}+\mathbf{E}$, and determine the rank $r_{k}$ such that $\mathbf{L}_{11}\in\mathbb{R}^{r_{k}\times r_{k}}$ and $\|\mathbf{E}\|_{F}\leq\varepsilon_{k}$.
@@
7:  $\mathbf{C}:=\mathbf{L}_{11}\mathbf{V}_{1}^{\top}$.
@@
[…]
Therefore, we also present the version of TT-ULV algorithm for prescribed accuracy in Algorithm 2 that can adaptively determine the TT-ranks. In this algorithm, if equal error weights are used, then $w_{k}=\frac{1}{\sqrt{d-1}}$. It is worth noting that in Step 5, we need to determine the smallest rank $r_{k}$ by comparing the truncation residual of the ULV decomposition with the given error tolerance $\varepsilon_{k}$. In the ULV desomposition $\mathbf{C}=\mathbf{U}_{1}\mathbf{L}_{11}\mathbf{V}_{1}^{\top}+\mathbf{E}$, the orthgonolity of $\mathbf{U}_{1}^{\top}\mathbf{E}=\mathbf{O}$ implies that $\|\mathbf{E}\|_{F}^{2}=\|\mathbf{C}\|_{F}^{2}-\|\mathbf{L}_{11}\|_{F}^{2}$. Thereby, we only need to compute the Frobenius norms of $\mathbf{C}$ and the row vectors of $\mathbf{L}_{11}$. Then the truncated rank $r_{k}$ can be determined as the smallest integer for which $\|\mathbf{L}_{11}\|_{F}^{2}\geq\|\mathbf{C}\|_{F}^{2}-\varepsilon_{k}^{2}$. If $\mathbf{C}$ is of size $m\times n$, then this process has a computational complexity of $O(mn)$, which is substantially lower than that of the preceding ULV decomposition.
As we can see, an essential requisite for the above error analysis for Algorithm 1 is that in each iteration the column-orthogonal matrix $\mathbf{U}_{1}$ used to generate the TT-core is orthogonal to the residual part $\mathbf{E}$ of the truncated ULV decomposition. Corollary 3.1 theoretically controls the accumulation of the truncated errors by a sharp bound.
However, if the right-orthogonal TT-cores are needed, the algorithm should adopt a right-to-left sweep, as given in Algorithm 4. Please refer to A, where we explain that this algorithm is not recommended for computing the right-orthogonal cores in the right-to-left sweep, especially for the objective of computing a fixed-precision TT decomposition. Interestingly, the URV case is precisely the opposite, which is more suitable to compute the TT format in a right-to-left sweep.
3.2 TT-URV case
Recall the URV decomposition (6)
$\mathbf{A}=\mathbf{U}_{1}\mathbf{R}_{11}\mathbf{V}_{1}^{\top}+(\mathbf{U}_{1}\mathbf{R}_{12}+\mathbf{U}_{2}\mathbf{R}_{22})\mathbf{V}_{2}^{\top},$
the matrix $\mathbf{V}_{1}$ is orthogonal to the residual part $(\mathbf{U}_{1}\mathbf{R}_{12}+\mathbf{U}_{2}\mathbf{R}_{22})\mathbf{V}_{2}^{\top}$. Wherefore, using the URV decomposition to compute the TT format is more suitable for the right-to-left sweep than for the left-to-right sweep.
We formulate the TT-URV algorithm 3 for fixed TT-ranks to replace Algorithm 4 in a right-to-left sweep. The results similar to Proposition 3.1 and Theorem 3.1 also hold for Algorithm 3.